# Tailored Forecasting from Short Time Series via Meta-learning

### v1

@@ Abstract
Machine learning (ML) models can be effective for forecasting the dynamics of unknown systems from time-series data, but they often require large amounts of data and struggle to generalize across systems with varying dynamics. Combined, these issues make forecasting from short time series particularly challenging. To address this problem, we introduce Meta-learning for Tailored Forecasting from Related Time Series (METAFORS), which uses related systems with longer time-series data to supplement limited data from the system of interest. By leveraging a library of models trained on related systems, METAFORS builds tailored models to forecast system evolution with limited data. Using a reservoir computing implementation and testing on simulated chaotic systems, we demonstrate METAFORS’ ability to predict both short-term dynamics and long-term statistics, even when test and related systems exhibit significantly different behaviors and the available data are scarce, highlighting its robustness and versatility in data-limited scenarios.
I Introduction
Accurate forecasting of dynamical systems is of practical importance across diverse domainsBrunton and Kutz (2019) including weather and climate scienceArcomano et al. (2020), neuroscienceWein et al. (2022); Pandarinath et al. (2018), epidemiologyRay et al. (2023), and financeSezer, Gudelek, and Ozbayoglu (2020). For many dynamical systems, however, we lack a sufficient knowledge-based model (i.e., mathematical or physical) to provide useful predictions. Data-driven approaches to infer a model of a system from its past evolution are highly valuable in these casesHan et al. (2021); Brunton, Proctor, and Kutz (2016). While powerful, these machine learning (ML) forecasters are typically data-intensive and brittleWang et al. (2021); Göring et al. (2024): they require lengthy training series and become inaccurate if the dynamics of the system to be predicted, i.e. the test system, differ even slightly from those seen in training. They thus struggle in many consequential scenarios. For example, public health officials responding to the initial outbreak of a virus may find that early data are insufficient to train an ML model for the spread of infection, and that one trained on data from another jurisdiction or previous outbreak does not generalize well. The motivation of this work is to facilitate useful forecasting when only a short time series is available from the system of interest. We focus on situations where no helpful knowledge-based model is available, but anticipate that our approach could also apply to hybrid architectures that integrate knowledge-based models with ML componentsPathak et al. (2018a); Arcomano et al. (2020).
The method we introduce amalgamates learned knowledge from ML forecasters trained on related, data-rich processes to improve performance on similarly related processes for which data are limited. As such, our work builds on prior studies in multi-task learning and meta-learning. These frameworks aim to reduce the data-intensity and brittleness of ML models by leveraging knowledge from tasks related to the one at hand.
Multi-task learning algorithms address data scarcity by training a single ML model on the union of all data from a set of training systems. No emphasis is placed on any particular training system or on the test system – the goal is to improve performance genericallyZhang and Yang (2022); Yang et al. (2020). In the context of time series forecasting, a multi-task learning approach would train a single model on a collection of time series from systems which each exhibit different dynamics. Multi-task learning may make the model less brittle to discrepancies between the dynamics of the training and test systems, but the training data that are most relevant to the behavior of the test system are often diluted with considerably less relevant dataZhang and Yang (2022). Thus, models trained by multi-task learning techniques may not achieve the desired performance for individual tasks.
Meta-learning, by contrast, assumes that sufficient data are available to train models for individual tasks. The goal of meta-learning, or ‘learning about learning’ is to generalize knowledge from these individual tasks to quickly adapt to new, unseen tasks. This process leverages prior learning to improve learning efficiency and performance specifically on individual future tasksBrazdil et al. (2022); Hospedales et al. (2022); Lemke, Budka, and Gabrys (2015); Pan and Yang (2010). Meta-learning has been used for a variety of purposesBrazdil et al. (2022); Lemke, Budka, and Gabrys (2015); Hospedales et al. (2022): hyperparameter optimizationFeurer, Springenberg, and Hutter (2015), algorithm selection and combinationLemke and Gabrys (2010); Talagala, Hyndman, and Athanasopoulos (2023), and few- or zero-shot learning. We are concerned mostly with this final goal – to tailor the trainable parameters of a model to a test system, given only a few data points from that system.
Figure 1: A Schematic of the METAFORS Meta-learning Method
(A) We train the forecaster separately on each of the available long training series, $\bm{L}_{i}$, to construct both a model representation of each corresponding dynamical system, and a cold-start vector at every time-step. (B) We then divide the long signals into short sub-signals, $\bm{s}_{ij}$, that are the same length as the available test signals and (C) train the signal mapper to map these short signals to cold-start vectors, $\bm{m}_{ij}$, corresponding to their start times, and appropriate model parameters, $\bm{\theta}_{i}$, for the forecaster, as in Sections IV.2.2, IV.2.3 and IV.2.4. (D) Given a test signal, $\bm{s}_{test}$, the signal mapper learns a suitable cold-start vector, $\bm{m}_{pi}$, and model parameters, $\bm{\theta}_{p}$. To make a prediction, we initialize the forecaster with the learned cold-start vector and drive the forecaster with the test signal to mitigate errors in the learned cold-start vector. Finally, we evolve the forecaster autonomously with the learned parameters.
@@ I Introduction
In this work, we construct a meta-learning approach for time series prediction that draws on a library of ML models with memory. By "memory" we mean that the output of the ML model depends not only on measurements of the current state of the system being forecasted, but also on measurements at previous times. This is the case for both ML architectures with intrinsic memory – eg. those based on recurrent neural networks, including long short-term memory models (LSTMs), gated recurrent units (GRUs), and reservoir computers (RCs), for which the memory is encoded in the network’s internal state – and ML architectures without intrinisic memory that explicitly include time-delayed measurements in their input – eg., next-generation reservoir computers and certain implementations of feedforward neural networks and kernel machines. ML models with memory are especially well-suited for time series prediction because they can extract information about unmeasured variables from the dynamics of measured variables. However, when we need to make forecasts from very short test signals, the model’s memory dependence creates an additional challenge: to predict accurately, the memory must be initialized consistently with the current state of the test systemGauthier et al. (2021); Zhang and Cornelius (2023); Lu, Hunt, and Ott (2018); Grigoryeva et al. (2024); Kemeth et al. (2021). Often, a substantial amount of data immediately preceding the start of a forecast is required for this purpose. In this case, if the available data are insufficient to initialize the ML model’s memory, forecasts will be either unobtainable or inaccurate, even if the model has been well-trained. In other words, memory-based ML models struggle to produce useful forecasts from a ‘cold start.’ We refer to a suitable initialization of a model’s internal state or memory as a ‘cold-start vector.’
In the scenarios we consider, the short time series data from the system of interest does not by itself provide enough information to accomplish the desired forecasting task. However, we assume that we are operating within the common situation where additional potentially relevant information is available that can be used alongside the short time series to enable accurate forecasting. Specifically, we assume that this additional information comes in the form of prior, more extensive measurements from other systems – and/or data from model simulations – whose dynamics we suspect are "related" to those of our system of interest.
@@
Meta-learning for Tailored Forecasting from Related Time Series (METAFORS), which we introduce here, allows memory-based ML forecasters to overcome the challenges of data-intensity, brittleness, and cold-starting. METAFORS exploits a collection of long training time series from related dynamical systems, using two levels of learning, to construct suitable models of new, similarly-related test systems for which data are limited. Two-level learning is a common feature in meta-learning approachesBrazdil et al. (2022).
@@
In our first learning level, we train separate forecasting models (Fig. 1A) for each long training series $\bm{L}_{i}(t)$. These models share the same base network architecture, and differ only in their trained model parameters, which we denote by a vector $\bm{\theta}_{i}$. Once trained, each model becomes a representation of its corresponding system’s dynamics. We also extract multiple short signals $\bm{s}_{ij}$ from each long signal, and during the training we record the model’s memory immediately prior to inputting $\bm{s}_{ij}$ as a cold-start vector $\bm{m}_{ij}$ (Fig. 1B). We store the short signals, along with their corresponding cold-start vectors and trained model parameters, in a ‘library’.
In the second learning level, another ML model, which we call the ‘signal mapper’, draws on this library to learn a direct mapping from short observation signals to appropriate forecaster parameters and cold-start vectors (Fig. 1C). The training data consists of inputs $\bm{s}_{ij}$ and corresponding target outputs $(\bm{m}_{ij},\bm{\theta}_{i})$. By leveraging this amalgamated knowledge, the signal mapper can then construct tailored forecasters for test signals with previously unseen or unknown dynamics (Fig. 1D). Importantly, METAFORS requires no explicit knowledge of the parameters or functional forms that govern the dynamics of either the library or test signals; it generalizes from the library models in an entirely data-driven manner.
Previous studies have shown that memory-based ML forecasters, when provided with explicit knowledge of either a relevant dynamical parameterPatel et al. (2021); Kong et al. (2021); Köglmayr and Räth (2024); Panahi and Lai (2024) or a contextual labelKong, Brewer, and Lai (2024) for each training and test series, can successfully forecast the long-term statistics, or climate, of test systems whose dynamics were not seen during training. These approaches are especially valuable in applications where we seek to predict a system’s behavior for a particular regime of a prescribed parameter—for example, projecting how the climate may evolve as atmospheric CO2 levels rise. By contrast, METAFORS tackles a related but distinct challenge: enabling rapid generalization to test systems where only short time series are available, and the relevant dynamical parameters remain unknown.
@@
The meta-learning framework – that is, learning to extract information from related tasks and utilize that information for the task at hand – facilitates this generalization in the absence of explicit context indicators. Several studies have explored meta-learning algorithms for zero-shot and few-shot prediction in applications beyond forecasting Finn, Abbeel, and Levine (2017); Yao et al. (2019); Raghu et al. (2020); Rusu et al. (2019); Wu et al. (2018). We are aware, however, of only a handful of works that directly address meta-learning for dynamical systems Canaday, Pomerance, and Girvan (2021); Kirchmeyer et al. (2022); Yin et al. (2021); Wang, Walters, and Yu (2022); Panahi et al. (2025). Among these approaches, METAFORS is important because: it does not rely on any specific neural network architectures; it both initializes and generalizes the forecaster model (and therefore readily accommodates forecasters with memory); it requires no awareness of dynamical parameters and little or no domain-knowledge; and it requires no re-training or optimization when presented with new test systems. METAFORS also demonstrates the versatility of meta-learning approaches to forecasting problems. It can forecast both climate and short-term evolution, generalize the forecaster to new test systems with qualitatively different behaviors, and learn to simultaneously represent dynamical systems of distinct functional forms.
@@
The way in which METAFORS deals with ML models with memory for forecasting is appealing in part because of existing and potential connections between information processing in artificial and biological neural networks. Recent workLu and Bassett (2020); Kim et al. (2020) has employed artificial neural network models with memory for time series forecasting to suggest biologically feasible learning paradigms that capture the flexibility of biological neural systems to learn different tasks simultaneously. However, these multi-task learning frameworks either manually separate the distributions of observed data from different tasksLu and Bassett (2020) or require a context-based control-parameterKim et al. (2020) to identify the task at hand. Here, we demonstrate that METAFORS can achieve the same flexibility without any explicit context-awareness and with considerable overlap in the distributions of the observed data.
To demonstrate METAFORS’ utility, we employ a similar reservoir computing implementation (Section IV.1) of memory-based ML forecasting models as used in the aforementioned papers that develop multi-task learning approaches for time series forecasting. Reservoir computers (RCs) have been shown to perform well for data-driven prediction and analysis of dynamical systemsSchrauwen, Verstraeten, and Campenhout (2007); Sun et al. (2024); Lukoševičius and Jaeger (2009), even those whose behavior is complex (e.g., dynamics on extended networksSrinivasan et al. (2022); Wikner et al. (2020)) or those that exhibit sensitivity to initial conditionsLukoševičius and Jaeger (2009); Tanaka et al. (2019); Lu et al. (2017); Pathak et al. (2018a); Krishnagopal et al. (2020); Pathak et al. (2018b); Patel et al. (2021); Wikner et al. (2020); Bollt (2021) (i.e., “chaotic” systems). Here, for simplicity, we use RCs as both the forecaster and the signal mapper. We emphasize, however, that in typical implementations, RCs also struggle with data-intensityLu, Hunt, and Ott (2018), brittlenessGöring et al. (2024); Yan et al. (2024), and cold-startingGauthier et al. (2021); Zhang and Cornelius (2023); Lu, Hunt, and Ott (2018); Grigoryeva et al. (2024).
In Section II, we show that METAFORS enables RC forecasting models to overcome each of these challenges. First, we demonstrate that METAFORS captures the long-term statistics, i.e. climate, of unseen test systems. Trained, for example, on just a few stationary trajectories from the chaotic regime of the logistic mapLorenz (1964), METAFORS can replicate the map’s dynamics across a large portion of its bifurcation diagram from short test signals with unknown dynamical parameters. Even when the library contains some ML models trained on logistic maps and others trained on Gauss iterated mapsHilborn (2000), METAFORS can construct the bifurcation diagrams of both – a difficult problem requiring it to implicitly determine the system type from a short test signal and build tailored forecasters in regions not covered in the library. In Section II.5, we show that METAFORS accurately forecasts the short-term evolution and climate of chaotic Lorenz-63 systemsLorenz (1963) with both fully-observed and partially-observed states. We highlight that when the available test signals are very short, the cold-start vectors learned by METAFORS are essential to accurate forecasting. Moreover, when the test and training systems are known to have the same underlying dynamics, METAFORS can be used to cold-start an RC forecaster with no more training data than is traditionally required. In Section III, we discuss the implications of our results and in Section IV we detail our implementation of RC forecasting models and the scope of METAFORS.
@@
Finally, although we illustrate our meta-learning method using machine learning (ML) based on reservoir computing, we emphasize that the general technique – combining short time series data from the system of interest with much more extensive data from related systems through a two-level meta-learning process – can be applied to other ML approaches with memory. These approaches will differ in the structural representations of both the ML model parameters for the library members and the encoded memory needed for cold-starting. Hence, the primary distinction in implementing METAFORS with different ML types is that the signal mapper must be trained on the different representations of model parameters and cold-start vectors. However, we do not anticipate this to present a fundamental challenge for implementing a METAFORS scheme. Furthermore, the method is also applicable to memoryless ML types, which could be relevant in cases where the time-series of the full state of the system is measured for both training and test signals. In the memoryless case, the cold-start vectors are not needed, and the METAFORS implementation simplifies significantly, as the signal mapper only has to map to the model parameters.
II Results
In this section, we first outline our reservoir computing implementation of METAFORS, leaving a more detailed description to Section IV. Then we use this implementation to show that METAFORS can generalize from a library of machine learning (ML) forecasting models to create tailored forecasts for test systems for which the dynamics are unknown and the available data are insufficient to train the forecaster directly. Our tests make use of data from three commonly studied nonlinear systems: the logistic mapLorenz (1964), the Gauss iterated mapHilborn (2000), and the Lorenz-63 equationsLorenz (1963).
@@ II.1 Implementation overview
The central component of a reservoir computer (RC) is a recurrent neural network, referred to as a reservoir. Each node in the reservoir network has an associated continuous-valued activation level, and the collection of all the node activations at a given time constitutes the state of the reservoir at that time, ${\bm{r}(t)}$. The reservoir state evolves over time in response to an external input signal, ${\bm{u}(t)}$, and is influenced, at every time step, by its own state at the previous time step via directed interactions between its nodes. This recurrence gives the reservoir ‘memory.’ We generate ‘outputs,’ ${\hat{\bm{u}}(t)}$, of the reservoir by forming linear combinations of the reservoir nodes’ activations, or, equivalently, applying a linear operator to the reservoir state vector. In this work, we train a forecaster RC for time series prediction by driving it with a training signal in ‘open-loop’ mode (Fig. 2A) and then choosing this linear operator, or ‘output layer’, such that the outputs at each time step closely match the target at the next time step, ${\hat{\bm{u}}(t)\approx\bm{u}(t+\Delta t)}$. We make this choice by simple linear regression and use a regularization parameter to prevent over-fitting. To make a prediction, we feed the RC’s output back as its input at each time step and it evolves as an autonomous dynamical system in ‘closed-loop’ mode (Fig. 2B). The RC’s output layer determines the dynamics of this autonomous system and its internal state, ${\bm{r}(t)}$, determines its position in the phase space of those dynamics. Typically, to initialize a forecast, we must synchronize the reservoir state to the test system by driving the RC in open-loop mode with a time series immediately preceding the start of the forecast. If the available time series is too short, the forecast will be inaccurate even if the RC’s output layer accurately represents the relevant dynamics. This is because the reservoir state contains memory about previous inputs, so it requires a sufficient ’synchronization time’ to properly initialize its memory.
@@
We build METAFORS’ library of long time series and associated trained ML models as follows. First, we train the forecaster on each available long signal as above. Through this process, we obtain, for each long signal, a well-trained output layer and well-synchronized reservoir states at every time step. The output layer constitutes the forecaster’s trained parameters, $\theta$, and its reservoir state at time $t$ constitutes a cold-start vector for a test signal starting at time $t$, ${\bm{m}(t)\equiv\bm{r}(t)}$. We train the signal mapper, which is also an RC in our implementation (but could easily be constructed by other ML schemes), to map each short signal extracted from the library to the reservoir state corresponding to its start and the output layer trained on the long signal from which it was extracted. In the prediction phase, the signal mapper learns, from a given test signal, an output layer for the forecaster and a cold-start vector corresponding to the start of the test signal. To mitigate errors in the learned cold-start vector, we then synchronize the forecaster to the test signal using this cold-start vector as the forecaster’s initial reservoir state. Finally, the forecaster evolves autonomously from the end of the test signal to generate a forecast.

### v2

@@ Abstract
Machine learning models can effectively forecast dynamical systems from time-series data, but they typically require large amounts of past data, making forecasting particularly challenging for systems with limited history. To overcome this, we introduce Meta-learning for Tailored Forecasting using Related Time Series (METAFORS), which generalizes knowledge across systems to enable forecasting in data-limited scenarios. By learning from a library of models trained on longer time series from potentially related systems, METAFORS builds and initializes a model tailored to short time-series data from the system of interest. Using a reservoir computing implementation and testing on simulated chaotic systems, we demonstrate that METAFORS can reliably predict both short-term dynamics and long-term statistics without requiring contextual labels. We see this even when test and related systems exhibit substantially different behaviors, highlighting METAFORS’ strengths in data-limited scenarios.
BrickRedIIntroduction
Forecasting of dynamical systems is crucial across fields such as weatherPrice et al. (2025) and climateArcomano et al. (2022) science, neuroscienceWein et al. (2022); De Matola and Miniussi (2025), epidemiologyRay et al. (2023), and financeSezer, Gudelek, and Ozbayoglu (2020). However, many systems lack accurate knowledge-based (e.g., mathematical or physical) models, necessitating data-driven approachesBrunton and Kutz (2019); Han et al. (2021); Brunton, Proctor, and Kutz (2016); Rudy et al. (2017) – either in standalone configurations or to augment insufficient models in hybrid configurationsPathak et al. (2018a); Arcomano et al. (2022). Standard machine learning (ML) forecasters, while powerful, are often ‘data intensive,’ requiring extensive training data to function effectively. They are also frequently ‘brittle,’ struggling to generalize across systems, even in cases where the systems’ dynamics are not very differentWang et al. (2021); Göring et al. (2024). For example, public health officials responding to a novel virus outbreak may find that early data are insufficient for training a new model, and models trained on previous outbreaks fail to generalize.
This work addresses the challenge of forecasting when only a short time series – insufficient to train a standalone model – is available from the system of interest. We assume access to longer time series from other related systems (and/or simulations) and that the long and short signals are both vector time series – sequences of measurement vectors over time. Our method, drawing on principles from multi-task learning and meta-learning, leverages knowledge from these related datasets to enable prediction or improve prediction accuracy while mitigating issues of data-intensity and brittleness.
While traditional multi-task learning aims to improve robustness by training a single model across multiple datasets, it can dilute system-specific informationZhang and Yang (2022); Yang et al. (2020). In contrast, meta-learning focuses on quickly adapting to new tasks by generalizing from task-specific modelsHospedales et al. (2022); Brazdil et al. (2022); Lemke, Budka, and Gabrys (2015); Pan and Yang (2010). Meta-learning has been applied in a range of settings, including hyperparameter optimizationFeurer, Springenberg, and Hutter (2015), algorithm selection and combinationLemke and Gabrys (2010); Talagala, Hyndman, and Athanasopoulos (2023), and few-shot learningFinn, Abbeel, and Levine (2017); Yao et al. (2019); Raghu et al. (2020); Rusu et al. (2019); Wu et al. (2018); Joshaghani et al. (2023); Canaday, Pomerance, and Girvan (2021); Kirchmeyer et al. (2022); Yin et al. (2021); Wang, Walters, and Yu (2022); Panahi et al. (2025); Oreshkin et al. (2021), which is our focus. Specifically, we wish to tailor the parameters of an ML model to forecast a dynamical system for which only a small amount of data is available.
@@
For individual time series prediction tasks, we focus on ML models with memory. By ‘memory,’ we mean that the output of the ML model depends not only on measurements of the current state of the system being forecast, but also on measurements at previous times. Such models include those with intrinsic memory encoded in an internal state – e.g., those based on recurrent neural networks, including long short-term memory models (LSTMs), gated recurrent units (GRUs), and reservoir computers (RCs) – and those without intrinsic memory that explicitly include time-delayed measurements in their input – e.g., next-generation reservoir computers (NGRCs) and certain feedforward neural networks and kernel machines. While memory-based ML models are well suited for time series prediction, they face an additional challenge when forecasting from short time series: accurate predictions rely on proper initialization of the model’s memory and often require substantial data from immediately before the start of the forecastGauthier et al. (2021); Zhang and Cornelius (2023); Lu, Hunt, and Ott (2018); Grigoryeva et al. (2024); Kemeth et al. (2021). We note that the data requirements for memory initialization and for forecaster training are different. It is possible to have enough data to initialize the model’s forecast without having enough to train its parameters. It is also possible to have enough data from the system of interest to train a good model while lacking sufficient data immediately before the start of the forecast, e.g., when using the model to forecast from a new short times series in cases where the underlying dynamics are expected to be the same as those seen in training. Without sufficient data for initialization, memory-based models struggle to produce useful forecasts from a ‘cold start,’ making forecasts inaccurate or unattainable. We refer to a suitable initialization of the model’s memory as a ‘cold-start vector.’
Figure 1: A Schematic of the METAFORS Meta-learning Method. (A) We train the forecaster separately on each of the available long training series, $\bm{L}_{i}$, to construct both a model representation of each corresponding dynamical system, and a cold-start vector at every time-step. (B) We then divide the long signals into short sub-signals, $\bm{s}_{ij}$, that are the same length as the short new system signals and (C) train the signal mapper to map these short signals to cold-start vectors, $\bm{m}_{ij}$, appropriate for their start times, and suitable model parameters, $\bm{\theta}_{i}$, for the forecaster. (D) Given a short new system signal, $\bm{s}_{new}$, the signal mapper learns a suitable cold-start vector, $\bm{m}_{n0}$, and model parameters, $\bm{\theta}_{n}$. To make a prediction, we initialize the forecaster with the learned cold-start vector and drive the forecaster with the short new system signal to mitigate errors in the learned cold-start vector. Finally, we evolve the forecaster autonomously with the learned parameters.
@@
To address the challenges of data limitations, brittleness, and cold-starting, we introduce Meta-learning for Tailored Forecasting using Related Time Series (METAFORS). METAFORS uses two-level learning (common in meta-learning schemesBrazdil et al. (2022)) to build tailored models for data-limited systems of interest by leveraging long time series from other systems suspected to be related.
@@
In our first learning level, we train separate forecasting models (Fig. 1A) for each long training series $\bm{L}_{i}(t)$. These models share the same base network architecture, and differ only in their trained model parameters, which we denote by a vector $\bm{\theta}_{i}$. We then extract multiple short signals $\bm{s}_{ij}$ from each long signal and, during the training, we record the state of the model’s memory – represented by the cold-start vector $\bm{m}_{ij}$ – immediately prior to inputting $\bm{s}_{ij}$ (Fig. 1B). These triplets (short signals, trained forecaster parameters, and cold-start vectors) form our training dataset for meta-learning.
In the second level (the meta-learning level) a ‘signal mapper’ ML model learns to map short observation signals in the training dataset to appropriate forecaster parameters and cold-start vectors (Fig. 1C). By leveraging knowledge across related systems, METAFORS generalizes to short signals from new systems without requiring explicit knowledge of governing equations or contextual labels (Fig. 1D).
@@
There are several aspects of METAFORS that differentiate it from related work. Previous studies have shown that memory-based ML forecasters, when provided with explicit knowledge of relevant dynamical parameters or contextual labels, can predict the long-term statistics (e.g., climate) of new systems with unseen dynamics, making them valuable for forecasting system behavior under specific parameter regimes, e.g., forecasting climate evolution in response to rising CO2 levelsPatel et al. (2021); Kong et al. (2021); Köglmayr and Räth (2024); Panahi and Lai (2024). METAFORS addresses a different challenge: rapid generalization to new systems with only a short time series and no explicit contextual indicators or knowledge of dynamical parameters. The meta-learning framework facilitates this context-free generalization. While meta-learning has been explored in forecasting and beyond Finn, Abbeel, and Levine (2017); Yao et al. (2019); Raghu et al. (2020); Rusu et al. (2019); Wu et al. (2018); Joshaghani et al. (2023); Talagala, Hyndman, and Athanasopoulos (2023), we are aware of only a few studies that focus on few-shot forecasting of dynamical systems Canaday, Pomerance, and Girvan (2021); Kirchmeyer et al. (2022); Yin et al. (2021); Wang, Walters, and Yu (2022); Panahi et al. (2025); Oreshkin et al. (2021). Among these approaches, METAFORS stands out because it does not rely on specific neural network architectures, both initializes and generalizes the forecaster model (accommodating memory-based models), requires no contextual information or domain knowledge, and needs no re-training when presented with new systems.
@@
The versatility that METAFORS confers to ML models with memory is appealing in part because of existing and potential connections between information processing in artificial and biological neural networks. METAFORS has overlap with Hopfield’s notion of associative memoryHopfield (1982); Kong, Brewer, and Lai (2024) in that it enables the recall of an entire pattern (full memory initialization) from just a partial input (short signal). It goes beyond associative memory by generalizing previously unseen partial inputs to appropriate memory initializations (cold-start vectors) and ML model parameters for that case. Other related recent workLu and Bassett (2020); Kim et al. (2021) has employed artificial neural network models with memory for time series forecasting to suggest biologically feasible learning paradigms that capture the flexibility of biological neural systems to learn different tasks simultaneously. These studies advance multi-task learning for time series forecasting by addressing task identification through either pre-processing that separates data distributions Lu and Bassett (2020) or by employing a contextual input Kim et al. (2021). METAFORS aims to achieve task flexibility without explicit context awareness or prior separation of tasks, enabling generalization even when data distributions overlap considerably.
@@
The general technique of METAFORS – combining a short ‘system signal’ from the system of interest with more abundant data from related systems through a two-level meta-learning process – can be applied broadly to ML approaches with memory. These approaches will differ in the structural representations of both the ML model parameters for the library members and the encoded memory needed for cold starting. However, we do not anticipate these varied representations to present a fundamental challenge for implementing a METAFORS scheme. The method is also applicable to memoryless ML types, for which its implementation simplifies considerably because the cold-start vectors are not needed and the signal mapper has to learn only model parameters.
To demonstrate METAFORS’ utility, we employ reservoir computers (RCs) as our memory-based ML forecasting models. For simplicity, we also use an RC for the signal mapper. RCs have been shown to perform well for data-driven prediction and analysis of dynamical systemsSchrauwen, Verstraeten, and Campenhout (2007); Sun et al. (2024); Lukoševičius and Jaeger (2009), even those whose behavior is complex (e.g., dynamics on extended networksSrinivasan et al. (2022)) or those that exhibit sensitivity to initial conditionsLukoševičius and Jaeger (2009); Tanaka et al. (2019); Lu et al. (2017); Pathak et al. (2018a); Krishnagopal et al. (2020); Pathak et al. (2018b); Patel et al. (2021); Wikner et al. (2020); Bollt (2021) (i.e., “chaotic” systems). We emphasize, however, that in typical implementations, RCs can also struggle with data-intensityLu, Hunt, and Ott (2018), brittlenessGöring et al. (2024); Yan et al. (2024), and cold startingGauthier et al. (2021); Zhang and Cornelius (2023); Lu, Hunt, and Ott (2018); Grigoryeva et al. (2024).
In this work, we show that METAFORS enables ML forecasting to overcome each of these challenges. Using simulated data from well-studied chaotic systems, we demonstrate that METAFORS can build tailored forecasters for systems with unseen and unknown dynamics using only short signals from these systems and no other contextual tags or domain-specific knowledge. We highlight that METAFORS captures both short-term evolution and long-term climate in several important scenarios: when the signals from the new system and the training systems exhibit substantially and qualitatively different dynamics; when the training signals originate from dynamical systems of distinct functional forms; and when the state of the underlying dynamical systems is only partially measured. Moreover, when the available system signals are very short, METAFORS’ cold-starting is essential to accurate forecasting, even when the new signal and the training signal(s) are known to have the same underlying dynamics.
BrickRedIIResults
We use toy chaotic systems as a controlled experimental setting to test METAFORS’ effectiveness. These systems – the logistic map, the Gauss iterated map, and the Lorenz-63 equations – are widely studied in nonlinear dynamics because they exhibit rich, complex behaviors despite being governed by relatively simple equations. Their well-characterized dynamics allow us to systematically generate ground truth data, and examine how system-relatedness, signal length, and library data coverage influence METAFORS’ ability to construct tailored forecasts. Moreover, these systems transition between qualitatively different behaviors with small changes in their dynamical parameters. This dynamical sensitivity demonstrates an important way in which a system of interest’s underlying dynamics may exacerbate the impact of ML model brittleness on forecast quality, and provides challenging conditions in which to test METAFORS’ generalization. In our experiments, we refer to the short system signals as ‘test signals,’ and we evaluate METAFORS against ground-truth data from these systems.
For the cases we consider, the data in both the long library signals and short test signals are discrete in time. This discreteness arises either because of sampling in the case of continuous-time systems, where we denote the sampling interval as ${\Delta t}$ (and take the ML prediction time step to also be ${\Delta t}$), or because the system itself evolves in discrete-time, as with the logistic map. For continuous-time systems, ${\Delta t}$ can be short compared to the system’s characteristic time scales (e.g., its Lyapunov time). By contrast, for discrete-time systems, the time step is typically on the scale of the system’s evolution dynamics.
The dynamical systems we study here all have attracting sets, or ‘attractors,’ in their state space, towards which trajectories starting within corresponding ‘basins of attraction’ evolve over time. A key challenge in forecasting from short time series is that insufficient sampling of the attractor’s dynamics limits the ability to approximate governing functions. More formally, learning a system’s dynamics corresponds to inferring a function – e.g., ${d\bm{x}/dt=\bm{F}_{\bm{p}}(\bm{x})}$ for ordinary differential equations or ${\bm{x}_{n+1}=\bm{M}_{\bm{p}}(\bm{x}_{n})}$ for discrete-time systems. METAFORS performs context-free learning of these functions using only observed state sequences ${\bm{x}(t)}$, without knowledge of system parameters ${\bm{p}}$, by leveraging information from systems with similar dynamics (i.e, similar functions ${\bm{F}}$ or ${\bm{M}}$) for which more data are available.
We first test METAFORS in an idealized setting where both the library and test signals come from the same type of system, the logistic map, and differ only in their parameters and initial conditions. Even in this simple case, where the dynamics follow a quadratic equation, standard machine learning methods that lack explicit knowledge of the governing equations or parameters struggle to generalize effectively. We then extend our analysis to a more challenging scenario, where the library and test signals are drawn from two distinct types of dynamical systems – the logistic map and the Gauss iterated map. Here, METAFORS must use only the test signal itself to construct a tailored forecasting model, without knowing which system generated the test signal nor any other contextual labels. Finally, using the Lorenz-63 equations, we show that METAFORS can successfully integrate information from both test and library signals to improve forecast accuracy, even when each signal contains only partial information of the state of the system that generated it.
We emphasize that every METAFORS library member consists of a long training signal and its corresponding ML model representation (trained parameters $\bm{\theta}$) only. To present results in pictorial form and to discuss context/parameter-aware baseline methods, we frequently refer to the dynamical parameters used to generate the long library signals and short test signals, e.g., the parameters of the logistic map or Lorenz equations. These dynamical parameters are not, however, included in METAFORS’ library and METAFORS can neither access them nor use them to make forecasts.
@@
Here, we provide an overview of our reservoir computing implementation of METAFORS and leave a more detailed description to the Methods section.
The central component of a reservoir computer (RC) is a recurrent neural network with fixed, randomly chosen links (see Section˜.1), referred to as the ‘reservoir’. Each node/neuron in the reservoir network has an associated continuous-valued activation level, and the collection of all node activations at a given time constitutes the state of the reservoir at that time, ${\bm{r}(t)}$. The reservoir state evolves over time in response to an external input signal, ${\bm{u}(t)}$, and is influenced, at every time step, by its own state at the previous time step via directed and weighted interactions between its nodes. This recurrence gives the reservoir ‘memory.’ We generate ‘outputs,’ ${\hat{\bm{u}}(t+\Delta t)}$, of the reservoir by forming linear combinations of the reservoir nodes’ activations, or, equivalently, applying a linear operator – called the ‘output layer’ – to the reservoir state vector. In this work, we train a forecaster RC for time series prediction by driving it with a training signal in ‘open-loop’ mode (Fig. 2A) and then choosing the output layer, such that the outputs at each time step closely match the target at the next time step, ${\hat{\bm{u}}(t+\Delta t)\approx\bm{u}(t+\Delta t)}$. We make this choice by simple linear regression, and use a regularization parameter to prevent over-fitting. To make a prediction, we feed the RC’s output back as its input at each time step and it evolves as an autonomous dynamical system in ‘closed-loop’ mode (Fig. 2B). The RC’s output layer determines the dynamics of this autonomous system and its internal state, ${\bm{r}(t)}$, determines its position in the state space of those dynamics. Typically, to initialize a forecast, we must synchronize the reservoir state to the test system by driving the RC in open-loop mode with a time series immediately preceding the start of the forecast. If the available time series is too short, the forecast will be inaccurate even if the RC’s output layer accurately represents the relevant dynamics. This is because the reservoir state contains memory about previous inputs, so it requires a sufficient ‘synchronization time’ to properly initialize its memory.
@@
We build METAFORS’ library of long time series and associated trained ML models as follows. First, using the same reservoir (e.g., with the same random realization of reservoir links), we train a forecaster on each available long signal. Through this process, we obtain, for each $\bm{L}_{i}$, using ${\bm{u}(t)=\bm{L}_{i}(t)}$, a well-trained output layer with trained model parameters $\bm{\theta}_{i}$ and well-synchronized reservoir states at every time step, which also serve as the cold-start vectors for short signals starting at time $t$, ${\bm{m}_{i}(t)=\bm{r}_{i}(t)}$. Next, we extract short signals from the long library signals and construct the training data for the signal mapper, comprising triplets of short signals $\bm{s}_{ij}$; corresponding RC model parameters $\bm{\theta}_{i}$; and cold-start vectors $\bm{m}_{ij}$ (state of the synchronized reservoir at the start of $\bm{s}_{ij}$). We then train the signal mapper, which is also an RC in our implementation (but could easily be constructed by other ML schemes), to associate the short signals with their corresponding RC model parameters and cold-start vectors. Once trained, the signal mapper is applied to a short new system signal, ${\bm{s}_{new}}$ to generate a tailored set of model parameters and a cold-start vector. To mitigate errors in the learned cold-start vector, we then run the forecaster in open-loop mode using ${\bm{u}(t)=\bm{s}_{new}(t)}$ with the cold-start vector as the forecaster’s initial reservoir state. Finally, the forecaster evolves autonomously from the end of the test signal to generate a forecast.